An air-filled capacitor consists of two parallel plates, each with an area of 7.60 cm2, separated by a distance of 2.10 mm. A 25.0-V potential difference is applied to these plates (a) Calculate the electric field between the plates kV/m (b) Calculate the surface charge density. nc/m2 (c) Calculate the capacitance. pF (d) Calculate the charge on each plate pc
Answer #1

A)

Electric field, E = V/d = 25/0.0021 = 11.90kV/m

B)

Density, s = E e0 = (11904.76 x 8.85 x 10^-12) = 105.36 nC/m^2

C)

Capacitance, C = Ae0/d = (0.00076 x 8.85 x 10^-12)/0.0021 = 3.20 pF

D)

Charge, Q = AEe0 = 0.00076 x 11904.76 x 8.85 x 10^-12 = 80.07 pC

Comment in case any doubt please rate my answer......
Answer #2

The 24 consonant sounds comprise six stops (plosives): p, b, t, d, k, g; the fricatives f, v, θ (as in thin), ð [eth] (as in then), s, z, [esh] (as in ship), Ʒ (as in pleasure), and h; two affricatives: t∫ (as in church) and dƷ (as the j in jam); the nasals m, n, ŋ (the sound that occurs at the end of words such as young); the lateral l; the postalveolar or retroflex r; and the semivowels j (often spelled y) and w. These remain fairly stable, but Inland Northern American differs from RP in two respects: (1) r following vowels is preserved in words such as door, flower, and harmony, whereas it is lost in RP; (2) t between vowels is voiced, so that metal and matter sound very much like British medal and madder, although the pronunciation of this t is softer and less aspirated, or breathy, than the d of British English.

answered by: Gamesome Goshawk